A simple pendulum complete one vibration in one sec. Find its length when = 10ms

1
Expert's answer
2022-05-22T14:38:28-0400
"T=2\pi\sqrt{l/g}"

"l=g(T/2\pi)^2\\
=10*(1/2\pi)^2=0.25\:\rm m"